3x-6=2x-5 show thw checking

Found 2 solutions by anamikakumari731@gmail.com, Alan3354:
Answer by anamikakumari731@gmail.com(1)

3x-6=2x-5
or 3x-2x =-5+6
or x=1 ans.

Answer by Alan3354(69443)

What does "thw" mean?